MrMudd183:

Can I get some assistance on factorials for Alg. III please? *I'll post an example of a question I'm stuck on below*

MrMudd183:

1 attachment
snowflake0531:

If you know what factorials are you have \[\frac{500!}{499!}=\frac{1 \times 2 \times3...\times 499 \times500 }{ 1 \times 2 \times 3... \times499}\] And if you know how to simplify the fraction, you can cross out all of the numbers from 1 to 499, because 1/1 is 1, 2/2 is 1 etc., so you're left with \(\frac{500}{1}=500\) do you get it?

MrMudd183:

Ahh makes sense, thank ya...and if it's a single digit such as 6! it'll be "6x5x4x3x2x1"?

snowflake0531:

yep

MrMudd183:

Great, thank ya!

snowflake0531:

yw~